Complete the sequence: 19, 5, 17, 21, 5, 14, 3, ____. Explain your answer.
Answer by askmemath(368)

Excluding the 5
19+2 = 21
17-3=14
The 5 just repeats itself every 3rd number.
The next number = 21+2 = 23.
5
14-3 = 11
The 3 however has me stumped.